# CADS: A Systematic Literature Review on the Challenges of Abstractive Dialogue Summarization

### v2

@@ 5.5 Human Evaluation
Human evaluation is often considered the gold standard for assessing the quality of a summary. It is typically performed through crowdsourcing annotators (e.g., Amazon Mechanical Turk) where they label samples. We identify established approaches for human evaluations on summary performance and annotator agreement.
@@ 6.1 Remarks on Challenges
Since 2023, NLP has seen a significant shift with the exploration of LLMs and optimized Transformer architecture components (e.g., ring attention (?) and multi-token generation (?)), with their application to dialogue summarization being explored later (?, ?, ?). Initial studies indicate that LLMs match or exceed the performance of task-specific models like DialogLED (?), even with their limitations (e.g., hallucinations, salience) (?). ? (?) show that encoder-decoder models may perform better against the speaker challenge, particularly regarding coreference resolution, and struggle with the structure challenge. LLMs handle the comprehension challenge but seem to be sub-optimal regarding robustness to language and speakers. These early observations state LLMs as a noteworthy alternative to more traditional techniques (e.g., dialogue-style pre-training (?), graph structures to represent speaker structures (?), and special losses tailored to determine salience (?)).
In the following, we discuss how the hurdles of the currently employed encoder-decoder models (identified in Section 3) align with challenges in NLP and discuss techniques introduced so far to mitigate these challenges, thereby deriving how they can aid dialogue summarization.
@@
Long distance handling and dependency parsing. Recent progress in LLMs such as GPT-4, Claude-3, and Phi-3 shows a significant extension of the processable context length to over 100k tokens, exceeding the context size of earlier models like LED (?), which handle up to 16k tokens. This increase in context allows for handling long dependencies without chunking, which is crucial for understanding complex dialogue structures, speaker dynamics, and content understanding. While traditional models could theoretically handle such lengthy contexts, practical limitations arise from the input-length-dependent quadratic computational costs associated with Transformer attention mechanisms (?). To address these issues, techniques like sparse attention (?), flash attention (training stage) (?) and ring attention (inference stage) (?), sharing weights across attention heads (?), and conditional computation to reduce the overall computational load (?) are proposed to manage large inputs.
Generalizability. Modifying pre-trained models to handle the salience and structure challenges independent of the conversation’s domain is essential for practical applications (e.g., customer service). Established generalization techniques involving prompting (?, ?), few-shot learning (?, ?), and meta-learning (?, ?) are incipient in dialogue summarization. Meanwhile, LLMs build on these techniques to leverage patterns observed during the training stage and boost their generalization capabilities (?). However, LLMs may still struggle with bias mitigation and ensuring broad applicability across various contexts, when ”encountered words are not present in the training corpus or in examples (?, ?, ?).
@@
Hallucination reduction. Avoiding hallucinated content in generated text is a major challenge in NLP tasks using LLMs, especially in culturally sensitive contexts (?, ?). Mitigation techniques include confidence penalty regularization (?, ?) to reduce overconfidence and enhance accuracy.
@@
  • We expect progress for the structure challenge with LLMs as the backbone model, but robustness issues may persist due to issues with generalization and processing long texts.
  • The comprehension challenge, especially in grasping implied meanings, remains largely unsolved and holds significant research potential. Research from related NLP fields such as sentiment analysis (?) suggest that current models still struggle to understand implied content. At the same time, research on retrieval-augmented generation (?) could enhance contextualization, helping models better grasp the context of conversations.
@@
  • The salience challenge currently builds on personalizing summaries and understanding long discussions, which both are gaining attention lately. Due to the novelty, the personalization of summaries requires more research on techniques and benchmarking datasets.
@@ 6.2 Trends in Datasets
Personalized Summarization. In contrast to the typically general summaries, personalized summaries tailored to the users’ needs are becoming increasingly popular (?, ?). Recent advancements have explored cross-attention and decoder self-attention to enhance role-specific information capturing (?). By integrating detailed personal attributes, models can better understand each participant’s background and motivations, leading to more targeted summaries. Dataset-wise, only a few, such as the Chinese CSDS (?), incorporate personalized summaries, whereas established English datasets typically provide only a single, general summary.
@@ 6.3 Trends in Evaluation
Evaluation metrics, vital for indicating performance and comparing new techniques, are currently mostly borrowed from related NLP tasks such as translation or text generation (Section 5). However, the effectiveness of the established metrics (e.g., ROUGE for dialogue summarization as identified in Table 1) is limited (?, ?), and while it can serve as a proxy (?) it provides insufficient insights into the true efficacy of new techniques. The also popular model-based metrics (e.g., BARTScore) seem unable to align well with human judgments (?) for dialogue summarization. Recent developments in NLP use LLM-based metrics such as GEMBA (?) and ICE (?) for evaluation (?), building on LLMs’ advanced text comprehension. This set of metrics thereby mimics human judgment and assesses common aspects (e.g., fluency, coverage, coherence) with continuous (?), Likert scale (?, ?), or probability scores (?). LLM-based metrics offer a promising direction for dialogue summarization evaluation due to their customizability, though this area remains largely unexplored.
@@ 7.1 Limitations of Evidence
Given our selection of papers on dialogue summarization, we found limitations in current works that may bias our findings. First, the crawled research studies in dialogue summarization predominantly focus on English-language dialogues, which renders their performance in multilingual or cross-cultural contexts unknown. While we frequently came across Chinese datasets, other languages are underrepresented. As a consequence, we focus our work on English dialogue summarization. Second, most datasets stem from specific sectors such as customer service or business meetings, which may not reflect the variety of dialogue types found in different environments. Lastly, the lack of detailed human evaluation reporting across research works, including inter-annotator agreement scores, further complicates the assessment of new techniques’ usefulness.
@@ 7.2 Limitations of Review Process
In our systematic review of dialogue summarization, we encountered methodological challenges that we had to circumvent. First, we focus our search on studies published in English and primarily involving English datasets. While this is a limiting factor, most research on the dialogue summarization task considers only English dialogues, making this a feasible approximation. However, we encourage others to assess the same issues we point out for other languages. Second, despite considering multiple databases, the timing and methods of our search may have missed studies, thus not fully capturing the dialogue summarization field. To mitigate this limitation, we updated our crawl several times throughout writing this article and retrieved works from two databases. We considered both peer-reviewed and non-peer-reviewed papers which risks including less credible studies or misinformation. To mitigate the potential weaknesses of non-peer-reviewed works, we use the ranking from semantic scholar and DBLP, which should account for the paper’s quality and evaluate the studies’ quality based on adherence to established reporting standards of top-tier conferences.
7.3 Conclusion
In this article, we reviewed works on Transformer-based abstractive dialogue summarization. We unified the existing definitions and concepts of dialogue summarization into a taxonomy (CADS) of the field’s six main challenge blocks: language, structure, comprehension, speaker, salience, and reliability-related factuality. We demonstrated how these challenges appear in dialogue summarization and discussed why they occur for established encoder-decoder-based summarization systems. We then grouped and highlighted techniques introduced since 2019 under the challenges they tackle. Despite advancements, we observed that the field is still in its infancy and offers ample opportunities for further research. We also listed datasets used in existing studies to illustrate the extent of data scarcity in dialogue summarization. Additionally, we evaluated common metrics and noted a strong reliance on the ROUGE metric, coupled with a lack of human evaluation reports, which raises doubts about the actual effectiveness of current techniques. In our discussion, we evaluated how current NLP techniques address these challenges and derived implications of using LLMs for the field. We thereby identify a lack of exploring such new models and borrowing effective techniques from related fields to overcome the limitations currently holding back progress. We conclude that while challenges like language may become less relevant, others, such as comprehension and factuality, still require more exploration. We recommend that future work should further discuss and adapt this taxonomy as new challenges, techniques, datasets, and evaluation measures emerge, and as current challenges may be solved naturally through technological innovation. The considered literature can be found in the accompanying repository  [^8] https://anonymous.4open.science/r/LitRev-DialogueSum/ which will be regularly updated.
@@ Acknowledgments
This work was supported by the Lower Saxony Ministry of Science and Culture and the VW Foundation. Frederic Kirstein was supported by the Mercedes-Benz AG Research and Development.
Glossary
LLM
Large Language Model
NLP
Natural Language Processing

### v3

@@ 5.5 Human Evaluation
Human evaluation is often considered the gold standard for assessing the quality of a summary. It is typically performed through crowdsourcing annotators (e.g., Amazon Mechanical Turk) who label samples. We identify established approaches for human evaluations on summary performance and annotator agreement.
@@ 6 Discussion
Throughout this work, we provide an overview of the current state of challenges, datasets, and evaluation. In Section 6.1, we analyze the emerging trends such as LLMs on mitigating the individual challenges, finding that our challenge taxonomy remains up to date. Section 6.2 discusses the increasing interest in datasets covering personalized summarization and realistic settings, while Section 6.3 covers recent research on improving evaluation metrics.
@@ 6.1 Remarks on Challenges
Since 2023, NLP has seen a significant shift with the exploration of LLMs and optimized Transformer architecture components (e.g., ring attention, ?, 2023, and multi-token generation, ?, 2024), with their application to dialogue summarization being explored later (?, ?, ?). Initial studies indicate that LLMs match or exceed the performance of task-specific models like DialogLED (?), even with their limitations (e.g., hallucinations, salience) (?). ? (?) show that encoder-decoder models may perform better against the speaker challenge, particularly regarding coreference resolution, and struggle with the structure challenge. LLMs handle the comprehension challenge but seem to be sub-optimal regarding robustness to language and speakers. These early observations state LLMs as a noteworthy alternative to more traditional techniques (e.g., dialogue-style pre-training, ?, 2022, graph structures to represent speaker structures, ?, 2022, and special losses tailored to determine salience, ?, 2023).
Following, we discuss how the hurdles of the currently employed encoder-decoder models (identified in Section 3) align with challenges in NLP and discuss techniques introduced to mitigate these challenges, thereby deriving how they can aid dialogue summarization.
@@
Long distance handling and dependency parsing. Recent progress in LLMs such as GPT-4, Claude-3, and Phi-3 shows a significant extension of the processable context length to over 100k tokens, exceeding the context size of earlier models like LED (?), which handle up to 16k tokens. This increase in context allows for handling long dependencies without chunking, which is crucial for understanding complex dialogue structures, speaker dynamics, and content understanding. While traditional models could theoretically handle such lengthy contexts, practical limitations arise from the input-length-dependent quadratic computational costs associated with Transformer attention mechanisms (?). To address these issues, techniques like sparse attention (?), flash attention during the training stage (?) and ring attention for inference (?), sharing weights across attention heads (?), and conditional computation to reduce the overall computational load (?) are proposed to manage large inputs.
Generalizability. Modifying pre-trained models to handle the salience and structure challenges independent of the conversation’s domain is essential for practical applications (e.g., customer service). Established generalization techniques involving prompting (?, ?), few-shot learning (?, ?), and meta-learning (?, ?) are incipient in dialogue summarization. Meanwhile, LLMs build on these techniques to leverage patterns observed during the training stage and boost their generalization capabilities (?). However, LLMs may still struggle with bias mitigation and ensuring broad applicability across various contexts when encountered words are not present in the training corpus or in examples (?, ?, ?).
@@
Hallucination reduction. Avoiding hallucinated content in generated text is a major challenge in NLP tasks using LLMs, especially in culturally sensitive contexts (?, ?). Mitigation techniques include confidence penalty regularization (?, ?) to reduce overconfidence and enhance accuracy and refinement methods to review and post-process a summary (?).
@@
  • We expect progress for the structure challenge using LLMs, but robustness issues may persist due to issues with generalization and processing long texts.
  • The comprehension challenge, especially in grasping implied meanings, remains largely unsolved and holds significant research potential. Research from related NLP fields such as sentiment analysis (?) suggests that current models still struggle to understand implied content. At the same time, research on retrieval-augmented generation (?) could enhance contextualization, helping models better grasp the context of conversations.
@@
  • The salience challenge builds on personalizing summaries and understanding long discussions, which both are gaining attention lately. Due to the novelty, the personalized summaries require more research on techniques and benchmarking datasets.
@@ 6.2 Trends in Datasets
Personalized Summarization. In contrast to the typically general summaries, personalized summaries tailored to the users’ needs are becoming increasingly popular (?, ?, ?). Recent advancements have explored cross-attention and decoder self-attention to enhance role-specific information capturing (?). By integrating detailed personal attributes, models can better understand each participant’s background and motivations, leading to more targeted summaries. Dataset-wise, only a few, such as the Chinese CSDS (?), incorporate personalized summaries, whereas established English datasets typically provide only a single, general summary.
@@ 6.3 Trends in Evaluation
Evaluation metrics, vital for indicating performance and comparing new techniques, are currently mostly borrowed from related NLP tasks such as translation or text generation (Section 5). However, the effectiveness of the established metrics (e.g., ROUGE for dialogue summarization as identified in Table 4) is limited (?, ?), and while it can serve as a proxy (?) it provides insufficient insights into the true efficacy of new techniques. The also popular model-based metrics (e.g., BARTScore) seem unable to align well with human judgments (?) for dialogue summarization. Recent developments in NLP use LLM-based metrics such as GEMBA (?) and ICE (?) for evaluation (?), building on LLMs’ advanced text comprehension. This set of metrics thereby mimics human judgment and assesses common aspects (e.g., fluency, coverage, coherence) with continuous (?), Likert scale (?, ?), or probability scores (?). LLM-based metrics offer a promising direction for dialogue summarization evaluation due to their customizability (?), though this area remains largely unexplored.
@@ 7 Final Considerations
In this article, we reviewed works on Transformer-based abstractive dialogue summarization. We unified the existing definitions and concepts of dialogue summarization into a taxonomy (CADS) of the field’s six main challenge blocks: language, structure, comprehension, speaker, salience, and reliability-related factuality. We demonstrated how these challenges appear in dialogue summarization and discussed why they occur for established encoder-decoder-based summarization systems. We then grouped and highlighted techniques introduced since 2019 under the challenges they tackle. Despite advancements, we observed that the field is still in its infancy and offers ample opportunities for further research. We also listed datasets used in existing studies to illustrate the extent of data scarcity in dialogue summarization. Additionally, we evaluated common metrics and noted a strong reliance on the ROUGE metric, coupled with a lack of human evaluation reports, which raises doubts about the actual effectiveness of current techniques. In our discussion, we evaluated how current NLP techniques address these challenges and derived implications of using LLMs for the field. We thereby identify a lack of exploring such new models and borrowing effective techniques from related fields to overcome the limitations currently holding back progress. We conclude that while challenges like language may become less relevant, others, such as comprehension and factuality, still require more exploration. We recommend that future work should further discuss and adapt this taxonomy as new challenges, techniques, datasets, and evaluation measures emerge and as current challenges may be solved naturally through technological innovation. The considered literature can be found in the accompanying repository [^9] https://github.com/FKIRSTE/LitRev-DialogueSum which will be regularly updated.
In the following, we discuss the discuss limitations of evidence (Section 7.1) and in our review process (Section 7.2).
@@ 7.1 Limitations of Evidence
Given our selection of papers on dialogue summarization, we found limitations in current works that may bias our findings. First, the crawled research studies in dialogue summarization predominantly focus on English-language dialogues, which renders their performance in multilingual or cross-cultural contexts unknown. While we frequently came across Chinese datasets, other languages are underrepresented. As a consequence, we focus our work on English dialogue summarization. Second, most datasets stem from specific sectors, such as customer service or business meetings, which may not reflect the various dialogue types in different environments. Lastly, the lack of detailed human evaluation reporting across research works, including inter-annotator agreement scores, further complicates the assessment of new techniques’ usefulness.
@@ 7.2 Limitations of Review Process
In our systematic review of dialogue summarization, we encountered methodological challenges that we had to mitigate. First, we focused our search on studies published in English and mainly using English datasets. While this is a limiting factor, most research on dialogue summarization considers only English dialogues, making this a feasible approximation. However, we encourage others to assess the issues we point out for other languages. Second, despite considering multiple databases, the timing and methods of our search may have missed studies, thus not fully capturing the dialogue summarization field. To mitigate this limitation, we updated our crawl several times throughout writing this article and retrieved works from two databases. We considered both peer-reviewed and non-peer-reviewed papers, which risks including less credible studies or misinformation. To mitigate the potential weaknesses of non-peer-reviewed works, we use the ranking from semantic scholar and DBLP, which should account for the paper’s quality and evaluate the studies’ quality based on adherence to established reporting standards of top-tier conferences.
@@ Acknowledgments
This work was supported by the Lower Saxony Ministry of Science and Culture and the VW Foundation. The Mercedes-Benz AG Research and Development supported Frederic Kirstein.